OpenStudy (anonymous):

A computer company gives a discount of 10 % on a new piece of software. If it makes a profit of 7%, what percent does the marked price exceed the cost price?

OpenStudy (kropot72):

Let marked price = M Discounted price = M - 0.1M = 0.9M Let cost price = C \[\frac{0.9M-C}{C}=0.07\] O.9M - C = 0.07C 1.07C = 0.9M \[\frac{M}{C}=\frac{1.07}{0.9}=1.1889\] The marked price exeeds the cost price by: (1.1889 - 1) * 100 = 18.89%